There are only two risky assets (stocks) A and B in the market.

Asset A: Mean = 20% Standard Deviation = 10%

Asset B: Mean = 10% Standard Deviation = 5%

Returns on Assets have zero correlation.

A.Assume that there is no risk-free asset.

(i)Plot (sketch) the efficiency frontier (the investment opportunity set).

(ii)What is the expected return and the standard deviation of the minimum-variance-portfolio?

(iii)An investor would like to construct a portfolio that has a standard deviation of 8%. Aconsultant suggest the following portfolio: ????=−0.393,????=1.393. Is this a goodsuggestion? Explain.
Answer #1
Expected Return of Stock A=R1 20%
Expected Return of Stock Y=R2 10%
S1=Standard Deviation of Stock A 10.00%
S2=Standard Deviation of Stock B 5.00%
Variance of Stock A=V1=(S1^2)         100 %%
Variance of Stock B=V2=(S2^2)      25.00 %%
Correlation (1,2) 0.00
Covariance(1,2)=Correlation*S1*S2 0.00 %%
w1=Investment in Stock A
w2=Investment in Stock B
Portfolio Return=Rp(Percentage)
w1*R1+w2*R2=w1*20+w2*10 ……..Equation (1)
Vp=Portfolio Variance=(w1^2)*V1+(w2^2)*V2+2*w1*w2*Covariance(1.2)
Vp=Portfolio Variance=(w1^2)*100+(w2^2)*25….....Equation(2)
Sp=Portfolio Standard Deviation=Square root of Variance=SQRT(Vp)
ALL POSSIBLE PORTFOLIOS
w1 w2 Rp=w1*20+w2*10 Vp(Using Equation (2) SP=Square Root(Vp)
Weight of Weight of
Stock A Stock B Portfolio Return(%) Portfolio Variance(%%) Portfolio Standard Deviation(%) Portfolio Return(%)
0 1 10 25.00 5.0% 10.0%
0.2 0.8 12 20.00 4.5% 12.0%
0.3 0.7 13 21.25 4.6% 13.0%
0.4 0.6 14 25.00 5.0% 14.0%
0.5 0.5 15 31.25 5.6% 15.0%
0.6 0.4 16 40.00 6.3% 16.0%
0.7 0.3 17 51.25 7.2% 17.0%
0.79 0.21 17.9 63.51 8.0% 17.9%
0.8 0.2 18 65.00 8.1% 18.0%
0.9 0.1 19 81.25 9.0% 19.0%
       1.00 0 20 100.00 10.0% 20.0%
      (0.39) 1.393 6.07 63.96 8.0% 6.1%
(ii) Minimum Variance Portfolio
Variance 20.00 %%
Expected Return 12.0%
Standard Deviation 4.5%
(iii) 8% Standard Deviation can be achieved by :
Weight of A=0.79, Weight of B=0.21
Expected Return of this portfolio 17.9%
If you use:
Weight of A=-0.393, Weight of B=1.393
Standard Deviation 8.0%
Expected Return of this portfolio 6.1%
This is not a good suggestion
Expected Return of this portfolio is much lower
Weight of A=0.79, Weight of B=0.21
Is a better solution for 8% standard Deviation

Efficiency Frontier: X-axis:Standard Deviation; Y axis-Return 25.0% 20.0% 15.0% 10.0% 5.0% 0.0% 0.0% 2.0% 4.0% 6.0% 8.0% 10.0